# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v3

@@ 4.1 Model Comparison
In particular, if checked more carefully, models with in-modal reconstruction decoders (e.g., CLIP-r and CLIP-pr compared with CLIP, and CLIP-pr compared with CLIP-p) generally show improved performance, as revealed by the “number of wins” in the table (last row), fairly accounting for the MLP-based downstream models only. This indicates enhanced informativeness of the learned representations for downstream parameter estimation. Overall, these results suggest that the inclusion of in-modal reconstruction decoders improves representation quality in most downstream applications.
We hypothesize that these performance gains stem from the model’s ability to retain both shared and modality-specific (non-shared) information. In the CLIP-pr variant, this is encouraged by jointly optimizing the contrastive loss, cross-modal prediction loss, and in-modal reconstruction loss. This architecture implicitly encourages the embeddings to retain complementary information from each modality.
@@
Beyond parameter estimation, we also evaluate the models on two additional tasks using 50,000 paired spectra selected from the validation split of the datasets used for contrastive training with decoders:
@@
These results are summarized in Table 2. We find that the models with both reconstruction and prediction decoders (CLIP-pr) yield improved performance on LRS $\rightarrow$ XP prediction, but slightly degrade the similarity score – an expected trade-off when the embeddings are trained to retain both shared and non-shared information. Nevertheless, their similarity scores remain significantly higher than the baseline similarity (0.0533) obtained from comparing the embeddings between modality-specific pre-trained models.
CLIP-split achieves the highest similarity score overall, even surpassing the baseline CLIP model, possibly aided by its lower embedding dimensionality, which tends to produce higher cosine similarities. It also delivers the best performance on XP $\rightarrow$ LRS spectrum prediction, demonstrating the model’s robustness across modalities.
In summary, models with both prediction and reconstruction decoders (CLIP-pr and CLIP-split) offer the best overall performance by balancing parameter-estimation accuracy, cross-modal predictability, and embedding similarity.
@@ 4.2.1 LAMOST LRS
For iron abundance [Fe/H], we benchmark our models against the DESI DR1 (DESI Collaboration et al., 2025; Koposov et al., 2025) and GALAH DR4 (Buder et al., 2025). For the GALAH comparison (33,411 stars), we retain stars with FLAG_FE_H = 0, SNR_PX_CCD2 $\geq$ 30, LAMOST SNRg $\geq$ 30 in both GALAH and LAMOST, and $T_{\rm eff}$ $\geq$ 4000 K. For the DESI comparison (119,335 stars), stars are selected with LAMOST SNRg $\geq$ 30, DESI SN_B $\geq$ 30, SUCCESS = True, and [Fe/H] $\geq-3.8$. Compared to DESI, the CLIP-split model demonstrates significantly better performance than the LAMOST official pipeline. The plateau in [Fe/H] around 2.5-2.5, which arises from the lack of metal-poor stars in the stellar library used by the LAMOST pipeline, is effectively addressed by our models, although the overall scatter remains comparable. When benchmarked against GALAH, the CLIP-split model achieves both lower bias and lower scatter, highlighting its competitive performance relative to physically motivated pipelines. We further investigate the impact of rest-frame correction by applying the predicted RVs to shift the spectra before feeding them into the [Fe/H] prediction models. This additional step yields predictions that are broadly consistent with those from uncorrected inputs, indicating that correcting for radial velocity is not critical—at least for the resolution and wavelength coverage of LAMOST LRS. The insensitivity of the MLP-based [Fe/H] predictions to small redshifts implies that the model has implicitly learned to accommodate these variations.
@@
Overall, our method compares favorably to previous work (Xiang et al., 2019; Li & Lin, 2023; Wang et al., 2023; Zhang et al., 2025; Zhao et al., 2025). Compared with the other data-driven methods listed, our approach generally requires fewer labeled training samples (typically<<90,000) and minimal hyperparameter tuning, as we adopt a unified architecture for all downstream models. Relative to physics-driven methods such as template fitting or forward modeling, our model does not require synthetic spectral templates or explicit physical modeling at inference time. This design, combined with diverse training data, enables applicability across a wide range of stellar types and delivers fast predictions once trained. While our method avoids physical modeling during inference, its effectiveness still depends on high-quality labels, all of which are ultimately derived from physics-based modeling approaches. These characteristics make our method both efficient and broadly applicable in practice. However, direct comparison with previous literature remains challenging due to differences in test set composition. For instance, our iron abundance test set extends down to [Fe/H] $\sim-4$, increasing the difficulty of achieving overall low scatter or high $R^{2}$.
4.2.2 Gaia XP
Figure 3 shows [Fe/H] predictions from the Gaia XP model (both pre-trained and CLIP-split models) compared with ground truth from the GALAH DR4 (Buder et al., 2025) and Gaia RVS catalog (Viswanathan et al., 2024). Our predictions are consistent across the entire iron abundance range, including the metal-poor regime down to about [Fe/H] =3.5=-3.5 or even 4.0-4.0, and outperform previous machine learning methods (Andrae et al., 2023), particularly at our ability to extend to low iron abundances. The overall scatter is below 0.08 dex for stars with [Fe/H] >2.0>-2.0, and below 0.18 dex for stars with [Fe/H] <2.0<-2.0. This performance even surpasses that achieved by traditional low-resolution spectroscopy.
Figure 4 highlights 135,370 extremely metal-poor (EMP) star candidates with iron abundances in the range $5<\mathrm{[Fe/H]}<-3$, identified by our XP branch of the CLIP-pr model. These stars exhibit a pronounced concentration toward the Galactic center, reminiscent of the “metal-poor heart of the Galaxy” reported by Rix et al. (2022), but now extending to significantly lower iron abundances than previously observed. A dedicated follow-up study based on this sample is currently underway, aiming to shed light on the earliest phases of the Milky Way’s chemical and structural evolution.
Performance metrics across various XP models are shown in Table 1. Again, CLIP-based models (CLIP, CLIP-pr, CLIP-split) are competitive compared with the earlier approaches, including Huang et al. (2024) and Li et al. (2024b). Notably, our test sets span a wide parameter range. For example, [Fe/H] extends down to around 4.0-4.0 and $T_{\rm eff}$ extends up to 13,500 K, further validating the robustness of our models.
@@ 4.2.2 Gaia XP
4.3 Spectral Retrieval and Prediction
4.3.1 Spectral Retrieval
@@
These capabilities suggest promising applications in data mining and search-based discovery. For instance, starting from a LAMOST spectrum of a rare type of stars, one could search for spectrally similar candidates in the Gaia XP database of over two billion stars. Such functionality could significantly enhance large-scale searches for rare or unusual stellar types.
4.3.2 Spectral Prediction
@@
We find that, for the majority of the test dataset, the model performs well in both directions (LRS $\rightarrow$ XP and XP $\rightarrow$ LRS), indicating that it effectively learns the mapping between these two spectroscopic modalities. Examples of such predictions are shown in Figure 5 as black curves.
However, for a subset of sources—particularly in the LRS $\rightarrow$ XP direction—prediction quality deteriorates. Apart from the potential effect of extinction poorly-learned by the trained model, this discrepancy may suggest that the source does not follow the behavior of a typical single star. For instance, it could be an unresolved binary or an otherwise anomalous object. These cases highlight an exciting future direction, using cross-modal prediction error as a basis for anomaly detection.
@@
5 Discussion
5.1 SBI Performance for [Fe/H] and Radial Velocity
In our experiments with SBI, we observed that applying NSF to projected embeddings led to degraded performance in the low-iron abundance regime ($\mathrm{[Fe/H]}\lesssim-2$). Similarly, radial velocity predictions from LAMOST LRS exhibited systematic underestimation of the absolute value at large radial velocities; a similar issue broadly discussed in machine learning-based estimators (Ting, 2024).
Interestingly, replacing NSF with MAF eliminated these issues. This suggests that MAF better models distributions near the parameter-space boundaries for these datasets. The contrasting performance between NSF and MAF highlights the importance of model choice in SBI, and motivates further investigation into the characteristics of these flows in astrophysical contexts.
5.2 Compression vs. Feature Learning
@@ 5.1 SBI Performance for [Fe/H] and Radial Velocity
This suggests that effective feature learning. rather than compression alone, is key to high downstream performance in this case. We note, however, that increasing the embedding dimension also increases the number of trainable parameters, which may partly explain the improved performance.
5.3 Interpretability of Parameter Estimation
@@
Recent techniques, such as sparse autoencoders (Cunningham et al., 2023), could enhance model explainability in future work. That said, our model achieves high precision and accuracy on held-out datasets, suggesting it is indeed learning physically meaningful representations. Further testing with carefully designed datasets will be necessary to validate this assumption and improve model interpretability.
5.4 SBI vs. MLP
Although we report MLP results for most tasks, we observe that SBI often outperforms MLP in terms of uncertainty (as measured by $\sigma$), while MLP yields better accuracy (as measured by $R^{2}$), as shown in Table 1.
This discrepancy arises from their differing training objectives; MLPs minimize the MSE, which aligns closely with $R^{2}$, whereas SBI focuses on modeling the full posterior distribution. In this work, we estimate the posterior median from SBI (rather than the mean), which is more robust to outliers.
Thus, for applications where uncertainty quantification is essential, SBI is the preferred choice. For fast and accurate point estimates, MLP is more suitable. Notably, our SBI models use only $\sim$0.1 million trainable parameters, compared to $\sim$1.4 million for the MLP models.
5.5 Training Sample Size and Dataset Configuration
We observe a positive correlation between training set size and model performance in downstream tasks. While our benchmark experiments use $\sim$100,000 stars per parameter, we note that performance may have not plateaued, indicating room for improvement with additional data.
Nevertheless, even with this moderate sample size, our models match or outperform state-of-the-art results in the literature (Section 4.2). To maximize data usage, we report metrics based on held-out test sets, while for plots generated using the downstream models, we combine the training and validation sets for training, with early stopping based on performance on the test set.
@@ 5.5 Training Sample Size and Dataset Configuration
5.6 Is Transformer Overkill?
In Table 4, we compare masked transformers (MT) and ordinary autoencoders (OAE) with equivalent numbers of trainable parameters and identical training epochs. For LAMOST LRS spectra, MT outperforms OAE. However, for Gaia XP spectra, OAE performs better.
These results suggest that transformers may be more advantageous for longer spectra (e.g., LRS with 1462 flux points), while offering limited benefits, or even unnecessary complexity, for shorter inputs like Gaia XP spectra (343 points). Another possible explanation is the difference in optimal training epochs for different architectures, as discussed in Różański & Ting (2025) and Różański et al. (2025). This warrants further exploration to better match model capacity to data complexity.
5.7 Fine-tuning Loss Weights
Not all non-shared information in the spectra is necessarily beneficial—some components, such as noise, may even hinder downstream performance. Moreover, the importance of shared versus non-shared information can vary across tasks; some may depend more on modality-specific features, while others benefit primarily from shared representations. Therefore, fine-tuning the weight terms in Equation 1, particularly the reconstruction loss weight, may be a promising direction for future work.
In this work, we fix both weights to 1, applying equal weighting to the reconstruction and cross-modal prediction losses. This choice shows that models incorporating reconstruction loss already perform competitively without explicit weight tuning, as evidenced by the number of top-performing metrics (i.e., “wins”) in Table 1 (last row)—e.g., CLIP-r vs. CLIP, and CLIP-pr/CLIP-split vs. CLIP-p. However, the magnitude of performance gains remains modest and somewhat task-dependent. Reweighting the losses may further increase the total number of improvements or yield more substantial gains on specific parameters. That said, the latter—optimizing for specific tasks—might come at the cost of generality, which is contrary to the foundational goal of building a model that performs robustly across diverse downstream tasks.
@@ 5.7 Fine-tuning Loss Weights
6 Summary
In this work, we develop a foundation model framework for stellar spectra that enables strong and efficient performance across multiple downstream tasks. Our approach integrates separate pre-trained models – each trained on a distinct spectroscopic modality (LAMOST LRS or Gaia XP) – and aligns them using CLIP-style contrastive learning. To further enhance the information capacity of the embeddings, we introduce decoder modules that increase the mutual information between the embeddings and input spectra and that enable translation (prediction) between different spectral types.

### v4

@@ V.1 Model Comparison
In particular, if checked more carefully, models with in-modal reconstruction decoders (e.g., CLIP-r and CLIP-pr compared with CLIP, and CLIP-pr compared with CLIP-p) generally show improved performance, as revealed by the “number of wins” in the table (last row), fairly accounting for the MLP-based downstream models only. This indicates enhanced informativeness of the learned representations for the estimation of downstream parameters. Overall, these results suggest that the inclusion of in-modal reconstruction decoders improves representation quality in most downstream applications.
We hypothesize that these performance gains come from the model’s ability to retain shared and modality-specific (non-shared) information. In the CLIP-pr variant, this is encouraged by jointly optimizing the contrastive loss, cross-modal prediction loss, and in-modal reconstruction loss. This architecture implicitly encourages the embeddings to retain complementary information from each modality.
@@
Beyond parameter estimation, we also evaluated the models on two additional tasks using 50,000 paired spectra selected from the validation split of the datasets used for contrastive training with decoders.
@@
These results are summarized in Table 2. We find that models with both reconstruction and prediction decoders (CLIP-pr) yield improved performance on LRS $\rightarrow$ XP prediction, but slightly degrade the similarity score, an expected trade-off when the embeddings are trained to retain both shared and non-shared information. Nevertheless, their similarity scores remain significantly higher than the baseline similarity (0.0533) obtained from comparing the embeddings between modality-specific pre-trained models.
CLIP-split achieves the highest similarity score overall, even surpassing the baseline CLIP model, possibly aided by its lower embedding dimensionality, which tends to produce higher cosine similarities. It also delivers the best performance on the prediction of the LRS spectrum of XP $\rightarrow$ LRS, demonstrating the robustness of the model in all modalities.
In summary, models with prediction and reconstruction decoders (CLIP-pr and CLIP-split) offer the best overall performance by balancing parameter-estimation accuracy, cross-modal predictability, and embedding similarity.
@@ V.2.1 LAMOST LRS
For iron abundance [Fe/H], we benchmark our models against the DESI DR1 and GALAH DR4. Compared to DESI, the CLIP-split model demonstrates significantly better performance than the LAMOST official pipeline. The plateau in [Fe/H] around 2.5-2.5, which arises from the lack of metal-poor stars in the stellar library used by the LAMOST pipeline, is effectively addressed by our models, although the overall scatter remains comparable. When benchmarked against GALAH, the CLIP-split model achieves both lower bias and lower scatter, highlighting its competitive performance relative to physically motivated pipelines. We further investigated the impact of rest-frame correction by applying the predicted RVs to shift the spectra before feeding them into the [Fe/H] prediction models. This additional step yields predictions that are broadly consistent with those from uncorrected inputs, indicating that correcting for radial velocity is not critical—at least for the resolution and wavelength coverage of LAMOST LRS. The insensitivity of the MLP-based [Fe/H] predictions to small redshifts implies that the model has implicitly learned to accommodate these variations.
@@
In general, our method compares favorably with previous work (2019ApJS..245...34X; 2023MNRAS.521.6354L; 2023ApJS..266...40W; 2025arXiv250602763Z; 2025ApJS..278...41Z). Compared with the other data-driven methods listed, our approach generally requires fewer labeled training samples (typically<<90,000) and minimal hyperparameter tuning, as we adopt a unified architecture for all downstream models. Relative to physics-driven methods such as template fitting or forward modeling, our model does not require synthetic spectral templates or explicit physical modeling at inference time. This design, combined with diverse training data, enables applicability across a wide range of stellar types and delivers fast predictions once trained. Although our method avoids physical modeling during inference, its effectiveness still depends on high-quality labels, all of which are ultimately derived from physics-based modeling approaches. These characteristics make our method both efficient and broadly applicable in practice. However, direct comparison with previous literature remains challenging due to differences in the composition of the test set. For example, our iron abundance test set extends to [Fe/H] $\sim-4$, increasing the difficulty of achieving a general low scatter or a high $R^{2}$.
V.2.2 Gaia XP
Figure 3 shows [Fe/H] predictions from the Gaia XP model (both pre-trained and CLIP-split models) compared with ground truth from the GALAH DR4 and Gaia RVS catalog. Our predictions are consistent across the entire iron abundance range, including the metal-poor regime down to about [Fe/H] =3.5=-3.5 or even 4.0-4.0, and outperform previous machine learning methods (2023ApJS..267....8A), particularly at our ability to extend to low iron abundances. The overall scatter is below 0.08 dex for stars with [Fe/H] >2.0>-2.0, and below 0.18 dex for stars with [Fe/H] <2.0<-2.0. This performance surpasses that achieved by traditional low-resolution spectroscopy.
Figure 4 highlights 135,370 extremely metal-poor (EMP) star candidates with iron abundances in the range $5<\mathrm{[Fe/H]}<-3$, identified by our XP branch of the CLIP-pr model. These stars exhibit a pronounced concentration toward the Galactic center, reminiscent of the “metal-poor heart of the Galaxy” reported by 2022ApJ...941...45R, but now extending to significantly lower iron abundances than previously observed. A dedicated follow-up study based on this sample is currently underway, aiming to shed light on the earliest phases of the Milky Way’s chemical and structural evolution.
Performance metrics in various XP models are shown in Table 1. Again, CLIP-based models (CLIP, CLIP-pr, CLIP-split) are competitive compared to earlier approaches, including 2024ApJ...974..192H and 2024ApJS..272....2L. Notably, our test sets span a wide parameter range. For example, [Fe/H] extends down to around 4.0-4.0 and $T_{\rm eff}$ extends up to 13,500 K, further validating the robustness of our models.
@@ V.2.2 Gaia XP
V.3 Spectral Retrieval and Prediction
V.3.1 Spectral Retrieval
@@
These capabilities suggest promising applications in data mining and search-based discovery. For instance, starting from a LAMOST spectrum of a rare type of star, one could search for spectrally similar candidates in the Gaia XP database of over two billion stars. Such functionality could significantly enhance large-scale searches for rare or unusual stellar types.
V.3.2 Spectral Prediction
@@
We find that for the majority of the test dataset, the model performs well in both directions (LRS $\rightarrow$ XP and XP $\rightarrow$ LRS), indicating that it effectively learns the mapping between these two spectroscopic modalities. Examples of such predictions are shown in Figure 5 as black curves.
However, for a subset of sources—particularly in the LRS $\rightarrow$ XP direction—prediction quality deteriorates. Apart from the potential effect of extinction, which may be poorly-captured by the trained model. This discrepancy may suggest that the source does not follow the behavior of a typical single star. For instance, it could be an unresolved binary or an otherwise anomalous object. These cases highlight an exciting future direction, using the cross-modal prediction error as a basis for anomaly detection.
@@
VI Discussion
VI.1 SBI Performance with NSF and MAF
In Table 1, we present the SBI results for several parameters ($T_{\rm eff}$, $\log g$, $v_{r}$, and asteroseismic parameters) that demonstrate the benefits of adopting SBI. For parameters showing relatively clear differences between NSF and MAF, we include results from both; for those exhibiting visually biased predictions in one of the two SBI methods, we conservatively report only the other. For parameters modeled by both methods (e.g., $T_{\rm eff}$), we observe better performance with NSF, likely due to its higher flexibility in representing non-linear conditional dependencies between spectra and stellar parameters.
In our additional experiments, we observed that applying NSF to projected embeddings led to degraded performance in the low-iron abundance regime ($\mathrm{[Fe/H]}\lesssim-1.5$). Similarly, LAMOST LRS radial velocity predictions exhibited a systematic underestimation of the absolute value at large radial velocities; a similar issue is widely discussed in machine learning-based estimators (2024arXiv241205806T). Interestingly, replacing NSF with MAF eliminated these issues.
These results suggest that NSF is more expressive than MAF, this increased expressiveness can adversely affect performance in our applications when modeling distributions near parameter-space boundaries. A likely explanation is that the spline transformations used in NSF, while highly flexible within their domain, become problematic in sparsely sampled boundary regions. Specifically, NSF employs rational–quadratic splines over a finite interval (default: $\pm 3\sigma$ in standardized space), with linear tail transformations beyond this range.
In astrophysical parameter spaces where extreme values ($\mathrm{Fe/H}]\lesssim-1.5$, $|v_{\mathrm{rad}}|\gtrsim 150\,\mathrm{km\,s^{-1}})$ represent several percent of the population, these linear tails systematically mis-model the true distribution shape. Furthermore, the flexibility of spline binning can lead to overfitting in regions with sparse training coverage, where bin allocation becomes poorly constrained.
In contrast, MAF’s simpler affine transformations naturally extend over the full unbounded support of the distribution with stable linear extrapolation, providing more robust behavior at distribution boundaries despite lower overall expressiveness. The contrasting performance between NSF and MAF highlights the importance of matching normalizing flow architecture to the characteristics of astrophysical parameter distributions, and motivates further investigation of boundary-aware flow designs for simulation-based inference in astronomy.
VI.2 Compression vs. Feature Learning
@@ VI.1 SBI Performance with NSF and MAF
This suggests that effective feature learning. rather than compression alone, is key to high downstream performance in this case. Mapping 343 flux points into a higher-dimensional latent space (e.g., 768) may allow the model to capture richer nonlinear correlations and disentangle latent physical factors (e.g., $T_{\rm eff}$, $\log g$, [Fe/H], extinction), forming expressive embeddings for downstream prediction. We note, however, that increasing the embedding dimension also increases the number of trainable parameters in the downstream MLP, which may partly explain the improved performance. These interpretations remain tentative and will be further examined in future work.
VI.3 Interpretability of Parameter Estimation
@@
Recent techniques, such as sparse autoencoders (2023arXiv230908600C), could improve the explainability of the model in future work. However, our model achieves high precision and accuracy on held-out datasets, suggesting that it is indeed learning physically meaningful representations. Further testing with carefully designed datasets will be necessary to validate this assumption and improve model interpretability.
VI.4 SBI vs. MLP
Although we report MLP results for most tasks, we observe that SBI tends to outperform MLP in terms of uncertainty (as measured by $\sigma$), while MLP yields comparable or better accuracy (as measured by $R^{2}$), as shown in Table 1.
This discrepancy may arise from their differing training objectives; MLPs minimize the MSE, which aligns closely with $R^{2}$, whereas SBI focuses on modeling the entire posterior distribution. In this work, we estimate the posterior median from SBI (rather than the mean), which is more robust to outliers. However, we note that differences in the adopted hyperparameters (see Appendix G) may also contribute to the discrepancy
Thus, for applications where uncertainty quantification is essential, SBI is the preferred choice. For fast and accurate point estimates, MLP is more suitable. In particular, our SBI models use only $\sim$0.1 million trainable parameters, compared to $\sim$1.4 million for the MLP models.
VI.5 Training Sample Size and Dataset Configuration
We observe a positive correlation between the size of the training set and the performance of the model in the downstream tasks. Although our benchmark experiments use $\sim$100,000 stars per parameter, we note that performance may not have plateaued, indicating room for improvement with additional data.
Nevertheless, even with this moderate sample size, our models match or outperform state-of-the-art results in the literature (Section V.2). To maximize data usage, we report metrics based on held-out test sets, while for plots generated using the downstream models, we combine the training and validation sets for training, with early stopping based on performance on the test set.
@@ VI.5 Training Sample Size and Dataset Configuration
VI.6 Is Transformer Overkill?
In Table 4, we compare MT and OAE with equivalent numbers of trainable parameters and identical training epochs. For LAMOST LRS spectra, MT outperforms OAE. However, for Gaia XP spectra, OAE performs better.
These results suggest that transformers may be more advantageous for longer spectra (e.g., LRS with 1462 flux points), while offering limited benefits, or even unnecessary complexity, for shorter inputs like Gaia XP spectra (343 points). Another possible explanation is the difference in optimal training epochs for different architectures, as discussed in 2025arXiv250318617R and 2025ApJ...980...66R. This warrants further exploration to better match model capacity to data complexity.
VI.7 Fine-tuning Loss Weights
Not all non-shared information in the spectra is necessarily beneficial—some components, such as noise, may even hinder downstream performance. Moreover, the importance of shared versus non-shared information can vary across tasks; some may depend more on modality-specific features, while others benefit primarily from shared representations. Therefore, fine-tuning the weight terms in Equation 1, particularly the loss weight of reconstruction, may be a promising direction for future work.
In this work, we fix both weights to 1, applying equal weighting to the reconstruction and cross-modal prediction losses. This choice shows that models incorporating reconstruction loss already perform competitively without explicit weight adjustment, as evidenced by the number of highest performing metrics (i.e. “wins”) in Table 1 (last row)—e.g., CLIP-r vs. CLIP, and CLIP-pr/CLIP-split vs. CLIP-p. However, the magnitude of performance gains remains modest and somewhat task-dependent. Reweighting the losses may further increase the total number of improvements or yield more substantial gains on specific parameters. However, the latter—optimizing for specific tasks—might come at the cost of generality, which is contrary to the fundamental goal of building a model that performs robustly across diverse downstream tasks.
@@ VI.7 Fine-tuning Loss Weights
VII Summary
In this work, we develop a foundation model framework for stellar spectra that enables strong and efficient performance across multiple downstream tasks. Our approach integrates separate pre-trained models, each trained on a distinct spectroscopic modality (LAMOST LRS or Gaia XP), and aligns them using CLIP-style contrastive learning. To further enhance the information capacity of the embeddings, we introduce decoder modules that increase the mutual information between the embeddings and input spectra and enable translation (prediction) between different spectral types.